# New measurements of the Lyman- forest continuum and effective optical depth with LyCAN and DESI Y1 data

### v1

@@ Abstract
We present the Lyman- Continuum Analysis Network (LyCAN), a Convolutional Neural Network that predicts the unabsorbed quasar continuum within the rest-frame wavelength range of based on the red side of the Lyman- emission line (). We developed synthetic spectra based on a Gaussian Mixture Model representation of Nonnegative Matrix Factorization (NMF) coefficients. These coefficients were derived from high-resolution, low-redshift () Hubble Space Telescope/Cosmic Origins Spectrograph quasar spectra. We supplemented this COS-based synthetic sample with an equal number of DESI Year 5 mock spectra. LyCAN performs extremely well on testing sets, achieving a median error in the forest region of 1.5% on the DESI mock sample, 2.0% on the COS-based synthetic sample, and 4.1% on the original COS spectra. LyCAN outperforms Principal Component Analysis (PCA)- and NMF-based prediction methods using the same training set by a factor of two or more. We predict the intrinsic continua of 83,635 DESI Year 1 spectra in the redshift range of and perform an absolute measurement of the evolution of the effective optical depth. This is the largest sample employed to measure the optical depth evolution to date. We fit a power-law of the form to our measurements and find and . Our results show particular agreement with high-resolution, ground-based observations around , indicating that LyCAN is able to predict the quasar continuum in the forest region with only spectral information outside the forest.
@@ 1 Introduction
While BOSS and eBOSS made great strides in constraining cosmological parameters, the nature of dark energy remains unknown. The Dark Energy Spectroscopic Instrument (DESI), the largest spectroscopic survey to date, aims to achieve the tightest constraints on the dark energy equation of state by measuring spectra of million galaxies and quasars over a five-year period (DESI Collaboration et al., 2016a, b). Roughly one million DESI targets are devoted to the high-redshift () quasars used to constrain cosmological parameters with the Ly forest. Each forest absorption feature provides a measurement of the matter-density distribution of the intergalactic medium (IGM) along the line of sight to the quasar. If the entire forest is visible, this can correspond to well over a hundred measurements of the matter-density field, relative to only one for quasars used as discrete tracers.
The Ly forest can also be used to probe properties of the IGM, such as its temperature, density, and ionization state. Measurements of the effective optical depth of the forest are one way to study the evolution of the mean density, temperature, photoionization rate, and the expansion history . The scatter of the optical depth also reveals information about the intensity of the ionizing background radiation produced by active galactic nuclei and star-forming galaxies (e.g. Becker et al., 2015; Kulkarni et al., 2019). While the Ly forest is readily accessible at with ground based observations, the optical depth is not straightforward to measure because it is difficult to determine the unabsorbed quasar continuum due to the forest absorption.
@@
where is the observed flux, and the mean expected flux is the product of the mean transmission and the unabsorbed continuum (du Mas desBourboux et al., 2020). These deltas are measured for each Ly forest pixel and then auto-correlated with each other or cross-correlated with quasars to achieve a 3D correlation function. Lacking knowledge of the true continuum, a two-parameter fit of the mean continuum in the forest region is performed on the observed flux. However, this introduces spurious correlations that complicate the analysis since the fitting process biases the mean of each delta toward zero for each line of sight. This biasing also results in the loss of information on large scales and projects out large-scale modes, which impacts the power spectrum in one dimension (e.g. Karaçaylı et al., 2020, 2022, 2024; Ravoux et al., 2023) and in three dimensions (e.g. Karim et al., 2023; de Belsunce et al., 2024). The standard approach is now to use a distortion matrix to accurately model and account for these spurious correlations, which is detailed in Bautista et al. (2017).
@@
Because a correlation exists between the continuum shape including broad emission lines on the red and blue sides of the Ly emission line (e.g. Pâris et al., 2011; Greig et al., 2017), other works have attempted to predict the true continuum using only longer wavelength information (redward of the Ly emission line) with Principal Component Analysis (PCA; e.g. Suzuki et al., 2005; Pâris et al., 2011; Lee et al., 2012; Davies et al., 2018). Lee et al. (2012) also corrected their continuum predictions using external constraints on the Ly forest mean flux, making their method dependent on the Ly forest region. In general, dimensionality reduction techniques like PCA or Nonnegative Matrix Factorization (NMF) are limited by their ability to learn linear representations of the data, while neural networks are able to learn more complex, nonlinear representations. We will show that PCA- and NMF-based prediction models do not work as well as a deep learning approach that has the ability to learn nonlinear relationships.
Several recent efforts have illustrated the potential of a deep learning approach. Liu & Bordoloi (2021) developed a feedforward neural network called iQNet designed to predict the quasar continuum in the Ly forest region using only information on the red side. However, their training set was entirely based on COS quasars and the resulting model is likely not optimized for DESI quasars, as for example the DESI quasars are generally much more luminous. Another effort from Sun et al. (2023) modeled the quasar continuum using an unsupervised probabilistic model. This model achieves promising results, although is not independent of the observed flux in the Ly forest region. The dependence on flux in the forest region also does not enable an independent measurement of the mean absorption.
@@ 2.1 HST/COS
We retrieved spectra from the Hubble Spectroscopic Legacy Archive (HSLA; Peeples et al., 2017), and then selected the subset with rest-frame wavelength coverage from and a median signal to noise ratio (SNR) of at least five per resolution element. We chose this minimum rest wavelength requirement to maximize our sample size, since very few quasars in the HSLA fully span the entire Ly forest region down to 1040 Å. Furthermore, some spectra suffer from a significant number of missing data pixels in this more narrow forest region. We removed QSOs that were missing more than 30% of the native pixels in the minimum forest region (), and then manually removed suspected Broad Absorption Line (BAL) quasars. This resulted in a sample of 48 QSOs. We transformed each quasar spectrum to units of relative flux by normalizing by the mean flux in the 2 Å interval centered around . Finally, we resampled these spectra to a rest wavelength grid from .
@@ 2.2 DESI
DESI was installed at the Mayall 4-m telescope at Kitt Peak National Observatory (DESI Collaboration et al., 2022) and is in the midst of the largest spectroscopic survey of the Universe. The aim of DESI is to measure spectra for million galaxies and quasars from the local Universe to (DESI Collaboration et al., 2016a) and study the nature of dark energy. The main cosmological tools are the BAO standard ruler to measure distances and constrain cosmological parameters, and redshift-space distortions to measure the growth of structure and test possible modifications to general relativity. DESI has 10 spectrographs and 5000 fibers that may be positioned across a 3 degree diameter field of view, which allows for efficient observations of 5000 targets simultaneously. The main DESI survey began in May 2021 upon the completion of commissioning and survey validation (DESI Collaboration et al., 2023a). An early data release included data from the survey validation period and the first two months of the main survey (DESI Collaboration et al., 2023b; Ramírez-Pérez et al., 2024). The DESI collaboration recently completed a series of papers that describe cosmological results from the first year of data collection (DESI Collaboration et al., 2024a, b, c). Additional key papers (DESI Collaboration et al., 2024d, e, f, g) based on Data Release 1 (DR1; DESI Collaboration et al., 2024h) are forthcoming.
@@ 2.3 DESI Mocks
We supplemented our synthetic dataset for training, testing, and validation with DESI mock spectra that were constructed for the main cosmological analysis (e.g. Farr et al., 2020; Herrera-Alcantar et al., 2023). The DESI mock spectra were generated with a software suite called quickquasars333https://github.com/desihub/desisim/blob/main/py/desisim/scripts/quickquasars.py. The quasar continua used by quickquasars are generated with the simqso444https://github.com/imcgreer/simqso library, which models the continuum as a broken power-law and is tuned to better reproduce the mean continuum observed in eBOSS DR16 (du Mas desBourboux et al., 2020) data via a Principal Component Analysis (PCA). These unabsorbed continua are multiplied by transmission files generated with the LyaCoLoRe program (Farr et al., 2020). quickquasars includes instrumental effects and astrophysical contaminants to produce “fully contaminated” mocks that include BALs, DLAs, high column density systems, and metal lines.
The final sample of mock spectra are representative of the DESI magnitude and redshift distribution, including with realistic noise and instrumental effects. For more details about the synthesis of the DESI mocks, see Herrera-Alcantar et al. (2023). For this work, we used the “fully contaminated” Year 5 (Y5) mock spectra, since these best represent the eventual quality of DESI data and already span the noise properties of Y1. The main distinction is that the Ly quasars will receive multiple observations over the course of the survey, and so the Y1 quasars will eventually have somewhat higher mean SNR. We used a realization of the mocks that contains the true continua in the forest region, in addition to the simulated (observed) forest. We conducted our analysis with a random sample of 40,000 mock spectra that exclude BALs and DLAs. These spectra fully cover the rest wavelength range of . This sample of mock spectra spans from .
@@ 3.2 Synthetic Spectra
After adding perturbations, we re-normalized each spectrum at 1280 Å. We then added random absorption features to the red side () of our perturbed continua to mimic metal absorption lines. Each absorption line is a Voigt profile with a Lorentzian full width at half maximum (FWHM) of 0.1 and a Gaussian FWHM of either 0.4 for the COS-like spectra or 1.0 for the DESI-like spectra. The Lorentzian amplitude was randomly selected between for the COS-like spectra or for the DESI-like spectra. These parameter values were chosen based on visual inspection of the COS and DESI Y1 spectra.
@@ 4.2 Network
We optimized the CNN architecture with a pseudo-random search in which we varied the number of convolutional layers, the kernel size and number of filters for these layers, the number of dense layers, and the number of nodes per dense layer. The number of convolutional layers ranged from one to four (including the input layer), and dense layers also ranged from one to four (not including the output layer). We trained and tested 10 different models for each possible combination (e.g. one convolutional layer and one hidden dense layer for the least complex models, up to four convolutional layers and four hidden dense layers for the most complex models). Each of these models was generated by randomly selecting the number of filters per convolutional layer and the number of nodes per hidden dense layer from a range of options. The number of filters was chosen to be a multiple of 16 between and the number of nodes was chosen to be a multiple of 32 between . The number of nodes in the output dense layer is always equal to the number of wavelength pixels in the full spectral range555Due to edge effects from resampling, the rest frame wavelength range actually spans from 1040.8-1599.2 Å (inclusive).. We separately trained these 160 configurations with a kernel size of three and a kernel size of six for a total of 320 unique neural network architectures.
We trained each of these unique architectures with the training subset of our data, and quantified each model’s performance on the testing subset. Among the testing subset, we evaluated the performance on the COS-based synthetic and DESI Y5 mock spectra with the Absolute Fractional Flux Error (AFFE), defined as
@@
We chose the architecture that minimized the median AFFE and range scores. This architecture features six hidden layers, three of which are convolutional layers and the remaining three are dense layers. The full architecture is shown in Table 1. LyCAN was compiled and trained using Keras666https://keras.io/ with the TensorFlow777https://www.tensorflow.org/ backend. We used the linear activation function for the output layer and the rectified linear unit (ReLU) activation function for all other layers, the mean squared error loss function, and the Adam optimizer. We trained for a maximum of 50 epochs and implemented early stopping criteria that monitored the validation loss starting at epoch 20 with a patience of five and a minimum delta of . The patience sets the number of epochs with negligible improvement after which training is stopped and optimal weights are restored, while the delta sets the minimum change in the validation loss to qualify as an improvement. There are a total of 169,851 trainable parameters (i.e. weights and biases) in our network. Lastly, the output continua were smoothed using a Gaussian filter with a standard deviation of three.
@@ 4.3 Performance and Comparisons
We created PCA- and NMF-based prediction models to compare the performance of these two other well-known methods to LyCAN. These models are based on the same training data as LyCAN, and are distinct from the previous NMF models developed to remove COS outliers and help generate COS-based synthetic spectra. We used scikit-learn888https://scikit-learn.org/stable/modules/generated/sklearn.decomposition.PCA.html to derive PCA components from the training set continua after z-score standardization. This scaled each spectrum so that its mean continuum is zero and standard deviation is one, which reduces possible biases due to luminosity. Once we computed the PCA continuum predictions, we reversed the standardization. For consistency, we used six components for both the PCA and NMF models. We fit the components to the red side of the training data () and then used the best-fit coefficients to extrapolate the continuum from the red side to the full spectral range, including the forest. Figure 9 shows a comparison of the PCA and NMF continuum predictions to LyCAN for four types of spectra. Neither technique works as well as LyCAN. In addition to their poorer performance, the components of the PCA and NMF models must be fit to smoothed continua in order to yield reasonable results. We therefore used the red side of the continua output by LyCAN as input to our PCA- and NMF-based models when predicting continua for DESI Y1 quasars.
@@
LyCAN outperforms these PCA- and NMF- based models by a factor of two or more. Figure 10 shows the distribution of AFFE values for different models and testing samples. The median AFFE values for different datasets and models are shown in Table 2. The first column shows the median AFFE values for these three models on the real COS quasars used to generate the synthetic data in the next two columns. LyCAN is able to predict the forest continua for the COS sample somewhat more accurately than the NMF and PCA models, although this sample is quite small. The next two columns of Table 2 list the median AFFEs of each model on the synthetic training and testing spectra, respectively, generated from the COS objects. Both of these subsets include synthetic spectra with DESI- and COS-like noise. LyCAN again outperforms the PCA and NMF models, achieving a median AFFE of 2.0%. We see nearly identical performance on the training versus testing subsets across all three models. Finally, the last two columns of Table 2 show the median AFFEs from each prediction model on the DESI Y5 mock training and testing sets. Again, LyCAN performs the best out of all the models, achieving a median AFFE of 1.5%, approximately two times better than the performance of the PCA model on the same dataset (3.3%) and more than six times better than the performance of the NMF model (9.3%). Interestingly, the NMF model performs more poorly than the PCA model in all cases except for the real COS sample. This is a surprising result, and may indicate that the NMF components are not as generalizable to more diverse datasets as the PCA components, whether that is due to the number of components used, the components themselves, or the standardization process. We also note that our PCA and NMF models have not been fully optimized, and more sophisticated PCA-based models in the literature that use more components, pixel weighting, and/or a projection matrix approach may offer better performance (e.g. Suzuki et al., 2005; Pâris et al., 2011; Lee et al., 2012; Davies et al., 2018).

### v2

@@ Abstract
We present the Lyman- Continuum Analysis Network (LyCAN), a Convolutional Neural Network that predicts the unabsorbed quasar continuum within the rest-frame wavelength range of based on the red side of the Lyman- emission line (). We developed synthetic spectra based on a Gaussian Mixture Model representation of Nonnegative Matrix Factorization (NMF) coefficients. These coefficients were derived from high-resolution, low-redshift () Hubble Space Telescope/Cosmic Origins Spectrograph quasar spectra. We supplemented this COS-based synthetic sample with an equal number of DESI Year 5 mock spectra. LyCAN performs extremely well on testing sets, achieving a median error in the forest region of 1.5% on the DESI mock sample, 2.0% on the COS-based synthetic sample, and 4.1% on the original COS spectra. LyCAN outperforms Principal Component Analysis (PCA)- and NMF-based prediction methods using the same training set by 40% or more. We predict the intrinsic continua of 83,635 DESI Year 1 spectra in the redshift range of and perform an absolute measurement of the evolution of the effective optical depth. This is the largest sample employed to measure the optical depth evolution to date. We fit a power-law of the form to our measurements and find and . Our results show particular agreement with high-resolution, ground-based observations around , indicating that LyCAN is able to predict the quasar continuum in the forest region with only spectral information outside the forest.
@@ 1 Introduction
While BOSS and eBOSS made great strides in constraining cosmological parameters, the nature of dark energy remains unknown. The Dark Energy Spectroscopic Instrument (DESI), the largest spectroscopic survey to date, aims to achieve the tightest constraints on the dark energy equation of state by measuring spectra of million galaxies and quasars over a five-year period (DESI Collaboration et al., 2016a, b). Roughly one million DESI targets are devoted to the high-redshift () quasars used to constrain cosmological parameters with the Ly forest. Each forest absorption feature provides a measurement of the matter-density distribution of the intergalactic medium (IGM) along the line of sight to the quasar. At the DESI resolution, this can correspond to well over a hundred independent absorption measurements per quasar if the entire forest is visible, whereas a quasar used as a discrete tracer only provides one measurement of the matter-density field.
The Ly forest can also be used to probe properties of the IGM, such as its temperature and ionization state (e.g. Schaye et al., 2000; Zaldarriaga et al., 2001; McDonald et al., 2001; Bolton et al., 2005; Meiksin, 2009; Becker et al., 2011, 2015; Walther et al., 2019; Khaire et al., 2019; Gaikwad et al., 2021). Measurements of the effective optical depth of the forest are one way to study these properties, and the scatter of the optical depth also reveals information about the intensity of the ionizing background radiation produced by active galactic nuclei and star-forming galaxies (e.g. Becker et al., 2015; Kulkarni et al., 2019). Other works have also studied the escape fraction of ionizing photons (e.g. Khaire et al., 2016) and the evolution of the background, including with the development of empirically-constrained models (e.g. Khaire & Srianand, 2019; Puchwein et al., 2019; Faucher-GiguÚre, 2020). McQuinn (2016) provided a comprehensive review on the evolution of the IGM. While the Ly forest is readily accessible at with ground based observations, the optical depth is not straightforward to measure because it is difficult to determine the unabsorbed quasar continuum due to the forest absorption.
@@
where is the observed flux, and the mean expected flux is the product of the mean transmission and the unabsorbed continuum (du Mas desBourboux et al., 2020). These deltas are measured for each Ly forest pixel and then auto-correlated with each other or cross-correlated with quasars to achieve a 3D correlation function. Lacking knowledge of the true continuum, a two-parameter fit of the mean continuum in the forest region is performed on the observed flux. However, this introduces spurious correlations that complicate the analysis since the fitting process biases the mean of each delta toward zero for each line of sight. This biasing also results in the loss of information on large scales and projects out large-scale modes, which impacts the power spectrum in one dimension (e.g. Karaçaylı et al., 2020, 2022, 2024; Ravoux et al., 2023) and in three dimensions (e.g. Karim et al., 2023; de Belsunce et al., 2024). The standard approach is now to use a distortion matrix to accurately model and account for these spurious correlations. Bautista et al. (2017) provide a detailed description of the distortion matrix.
@@
Because a correlation exists between the continuum shape including broad emission lines on the red and blue sides of the Ly emission line (e.g. Pùris et al., 2011; Greig et al., 2017), other works have attempted to predict the true continuum using only longer wavelength information (redward of the Ly emission line) with Principal Component Analysis (PCA; e.g. Suzuki et al., 2005; Pùris et al., 2011; Lee et al., 2012; Davies et al., 2018). Lee et al. (2012) also corrected their continuum predictions using external constraints on the Ly forest mean flux, making their method dependent on the Ly forest region. In general, dimensionality reduction techniques like PCA or Nonnegative Matrix Factorization (NMF) are limited by their ability to learn linear representations of the data, while neural networks are able to learn more complex, nonlinear representations. We will show that PCA- and NMF-based prediction models do not work as well as a deep learning approach that has the ability to learn nonlinear relationships.
Several recent efforts have illustrated the potential of deep learning approaches. Some have used such approaches to infer properties of the IGM (e.g. Huang et al., 2021; Nayak et al., 2023; Nasir et al., 2024) or fit the absorption lines in the forest with Voigt profiles (Jalan et al., 2024). Additionally, deep learning can be used to predict the unabsorbed continuum in the forest region. Liu & Bordoloi (2021) developed a feedforward neural network called iQNet designed to predict the quasar continuum in the Ly forest region using only information on the red side. However, their training set was entirely based on COS quasars and the resulting model is likely not optimized for DESI quasars, as for example the DESI quasars are generally much more luminous. Another effort from Sun et al. (2023) modeled the quasar continuum using an unsupervised probabilistic model. This model achieves promising results, although is not independent of the observed flux in the Ly forest region. The dependence on flux in the forest region also does not enable an independent measurement of the mean absorption.
@@ 2.1 HST/COS
We retrieved spectra from the Hubble Spectroscopic Legacy Archive (HSLA; Peeples et al., 2017), and then selected the subset with rest-frame wavelength coverage from and a median signal to noise ratio (SNR) of at least five per resolution element. We chose this minimum rest wavelength requirement to maximize our sample size, since very few quasars in the HSLA fully span the entire Ly forest region down to 1040âĂ . Furthermore, some spectra suffer from a significant number of missing data pixels in this more narrow forest region. We removed QSOs that were missing more than 30% of the native pixels in the minimum forest region (), and then manually removed suspected Broad Absorption Line (BAL) quasars. This resulted in a sample of 48 QSOs. Finally, we transformed each quasar spectrum to units of relative flux by normalizing by the mean flux in the 2âĂ â interval centered around .
@@ 2.2 DESI
DESI was installed at the Mayall 4-m telescope at Kitt Peak National Observatory (DESI Collaboration et al., 2022) and is in the midst of the largest spectroscopic survey of the Universe. The aim of DESI is to measure spectra for million galaxies and quasars from the local Universe to (DESI Collaboration et al., 2016a) and study the nature of dark energy. The main cosmological tools are the BAO standard ruler to measure distances and constrain cosmological parameters, and redshift-space distortions to measure the growth of structure and test possible modifications to general relativity. DESI has 10 spectrographs and 5000 fibers that may be positioned across a 3 degree diameter field of view, which allows for efficient observations of 5000 targets simultaneously. The main DESI survey began in May 2021 upon the completion of commissioning and survey validation (DESI Collaboration et al., 2023a). An early data release included data from the survey validation period and the first two months of the main survey (DESI Collaboration et al., 2023b; RamĂrez-PĂ©rez et al., 2024). The DESI collaboration recently completed a series of papers that describe cosmological results from the first year of data collection (DESI Collaboration et al., 2024a, b, c). Additional key papers (DESI Collaboration et al., 2024d, e, f, g) based on Data Release 1 (DR1; DESI Collaboration et al., 2024h) are forthcoming.
@@ 2.3 DESI Mocks
We supplemented our synthetic dataset for training, testing, and validation with DESI mock spectra that were constructed for the main cosmological analysis (e.g. Farr et al., 2020; Herrera-Alcantar et al., 2023). The DESI mock spectra were generated with a software suite called quickquasars333https://github.com/desihub/desisim/blob/main/py/desisim/scripts/quickquasars.py. The quasar continua used by quickquasars are generated with the simqso444https://github.com/imcgreer/simqso library, which models the continuum as a broken power-law and is tuned to better reproduce the mean continuum observed in eBOSS DR16 (du Mas desBourboux et al., 2020) data via a Principal Component Analysis (PCA). These unabsorbed continua are multiplied by transmission files generated with the LyaCoLoRe program (Farr et al., 2020). quickquasars includes instrumental effects and astrophysical contaminants to produce âfully contaminatedâ mocks that include BALs, DLAs, high column density systems, and metal lines.
The final sample of mock spectra are representative of the DESI magnitude and redshift distribution, including with realistic noise and instrumental effects. For more details about the synthesis of the DESI mocks, see Herrera-Alcantar et al. (2023). For this work, we used the âfully contaminatedâ Year 5 (Y5) mock spectra, since these best represent the eventual quality of DESI data and already span the noise properties of Y1. The main distinction is that the Ly quasars will receive multiple observations over the course of the survey, and so the Y1 quasars will eventually have somewhat higher mean SNR. We used a realization of the mocks that contains the true continua in the forest region, in addition to the simulated (observed) forest. We conducted our analysis with a random sample of 40,000 mock spectra that exclude BALs and DLAs. These spectra fully cover the rest wavelength range of . This sample of mock spectra spans from .
@@ 3.2 Synthetic Spectra
After adding perturbations, we re-normalized each spectrum at 1280âĂ . We then added random absorption features to the red side () of our perturbed continua to mimic metal absorption lines. Each absorption line is a Voigt profile with a Lorentzian full width at half maximum (FWHM) of 0.1 and a Gaussian FWHM of either 0.4 for the COS-like spectra or 1.0 for the DESI-like spectra. The Lorentzian amplitude was randomly selected between for the COS-like spectra or for the DESI-like spectra. The range of absorption features and the parameter values were chosen based on visual inspection of the COS and DESI Y1 spectra.
@@ 4.2 Network
We optimized the CNN architecture with a pseudo-random search in which we varied the number of convolutional layers, the kernel size and number of filters for these layers, the number of dense layers, and the number of nodes per dense layer. The number of convolutional layers ranged from one to four (including the input layer), and dense layers also ranged from one to four (not including the output layer). We trained and tested 10 different models for each possible combination (e.g. one convolutional layer and one hidden dense layer for the least complex models, up to four convolutional layers and four hidden dense layers for the most complex models). Each of these models was generated by randomly selecting the number of filters per convolutional layer and the number of nodes per hidden dense layer from a range of options. The number of filters was chosen to be a multiple of 16 between and the number of nodes was chosen to be a multiple of 32 between . The number of nodes in the output dense layer is always equal to the number of wavelength pixels in the full spectral range555Due to edge effects from resampling, the rest frame wavelength range actually spans from 1040.8-1599.2Â Ă Â (inclusive).. We separately trained these 160 configurations with a kernel size of three and a kernel size of six for a total of 320 unique neural network architectures.
We trained each of these unique architectures with the training subset of our data, and quantified each modelâs performance on the testing subset. Among the testing subset, we evaluated the performance on the COS-based synthetic and DESI Y5 mock spectra with the Absolute Fractional Flux Error (AFFE), defined as
@@
We chose the architecture that minimized the median AFFE and range scores. This architecture features six hidden layers, three of which are convolutional layers and the remaining three are dense layers. The full architecture is shown in Table 1. LyCAN was compiled and trained using Keras666https://keras.io/ with the TensorFlow777https://www.tensorflow.org/ backend. We used the linear activation function for the output layer and the rectified linear unit (ReLU; Krizhevsky et al., 2012) activation function for all other layers, the mean squared error loss function, and the Adam optimizer (Kingma & Ba, 2017). We trained for a maximum of 50 epochs and implemented early stopping criteria that monitored the validation loss starting at epoch 20 with a patience of five and a minimum delta of . The patience sets the number of epochs with negligible improvement after which training is stopped and optimal weights are restored, while the delta sets the minimum change in the validation loss to qualify as an improvement. There are a total of 169,851 trainable parameters (i.e. weights and biases) in our network. Lastly, the output continua were smoothed using a Gaussian filter with a standard deviation of three in order to reduce the noise in the predictions and minimize the AFFEs.
@@ 4.3 Performance and Comparisons
We created PCA- and NMF-based prediction models to compare the performance of these two other well-known methods to LyCAN. These models are based on the same training data as LyCAN, and are distinct from the previous NMF models developed to remove COS outliers and help generate COS-based synthetic spectra. We used scikit-learn888https://scikit-learn.org/stable/modules/generated/sklearn.decomposition.PCA.html to derive PCA components from the training set continua after z-score standardization. This scaled each spectrum so that its mean continuum is zero and standard deviation is one, which reduces possible biases due to luminosity. Once we computed the PCA continuum predictions, we reversed the standardization. We optimized the PCA and NMF models by testing models with up to components. For each model, we fit the components to the red side of the training data () and then used the best-fit coefficients to extrapolate the continuum from the red side to the full spectral range, including the forest. We used the median AFFE on the DESI mock testing set as a performance metric to find the optimal models. For both PCA and NMF, we found that four components minimized this metric, and we therefore adopted these as the optimal models.
Figure 9 shows a comparison of the PCA and NMF continuum predictions to LyCAN for four types of spectra. Neither technique works as well as LyCAN. In addition to their poorer performance, the components of the PCA and NMF models must be fit to smoothed continua in order to yield reasonable results. We therefore used the red side of the continua output by LyCAN as input to our PCA- and NMF-based models when predicting continua for DESI Y1 quasars.
@@
LyCAN outperforms these PCA- and NMF- based models. Figure 10 shows the distribution of AFFE values for different models and testing samples. The median AFFE values for different datasets and models are shown in Table 2. The first column shows the median AFFE values for these three models on the real COS quasars used to generate the synthetic data in the next two columns. LyCAN is able to predict the forest continua for the COS sample more accurately than the NMF and PCA models, although this sample is quite small. The next two columns of Table 2 list the median AFFEs of each model on the synthetic training and testing spectra, respectively, generated from the COS objects. Both of these subsets include synthetic spectra with DESI- and COS-like noise. LyCAN again outperforms the PCA and NMF models, achieving a median AFFE of 2.0%. We see nearly identical performance on the training versus testing subsets across all three models. Finally, the last two columns of Table 2 show the median AFFEs from each prediction model on the DESI Y5 mock training and testing sets. Again, LyCAN performs the best out of all the models, achieving a median AFFE of 1.5%, approximately 40% better than the performance of the PCA model on the same dataset (2.6%) and almost five times better than the performance of the NMF model (7.4%). Interestingly, the NMF model performs more poorly than the PCA model in all cases except for the real COS sample. This is a surprising result, and may indicate that the NMF components are not as generalizable to more diverse datasets as the PCA components, whether that is due to the number of components used, the components themselves, or the standardization process. We also note that while the number of components has been optimized for our PCA and NMF models, more sophisticated PCA-based models in the literature that use pixel weighting and/or a projection matrix approach may offer better performance (e.g. Suzuki et al., 2005; Pùris et al., 2011; Lee et al., 2012; Davies et al., 2018).